OpenStudy (anonymous):

meadals an fan!!!!!!!!!!!!!!!!!!!!!!!!![3.02]What is the greatest common factor of fifteen and thirty-five? 3 5 7 fifteen

OpenStudy (anonymous):

5

OpenStudy (anonymous):

factors of 15 are 1, 3, 5, and 15 factors of 35 are 1, 5, 7, and 35

OpenStudy (anonymous):

3.05]When one U.S. Dollar is worth one point four eight Australian Dollars, which of the following could be used as a conversion factor? two hundred fifty thirty-sevenths one and twelve twenty-fifths one and twenty-nine fiftieths two and thirteen twenty-fifths @juliannescriv

OpenStudy (anonymous):

one and twelve twenty0fifths

OpenStudy (anonymous):

b

OpenStudy (anonymous):

1.48=1+0.48=1+48/100=1+12/25

OpenStudy (anonymous):

Which has 8 as a factor? 4 twelve twenty-four thirty-eight @juliannescriv

OpenStudy (anonymous):

24

OpenStudy (anonymous):

8 doesnt go into any of those numbers except 24

OpenStudy (anonymous):

[3.02]Which of the following is in simplest form? seven thirty-fifths eighteen fifty-fourths thirty forty-fifths nineteen twenty-sevenths @juliannescriv

OpenStudy (anonymous):

d

OpenStudy (anonymous):

Derek distributed twenty-eight pieces of apple to his friends. If each apple slice was one-eighth of an apple, how many apples did he distribute to his friends? three and one-half four and one-half three and three-fourths four and one-fourth @juliannescriv

OpenStudy (anonymous):

3 1/2

OpenStudy (anonymous):

3.01]Which of the following is true? fifteen is a prime number because it has a factor of 5. eleven is a composite number because one and eleven are its only factors. 4 is a prime number because it is only divisible by one and itself. 9 is a composite number because it is divisible by 3. @juliannescriv

OpenStudy (anonymous):

d

OpenStudy (anonymous):

3.01]Which of the following helps you to determine that three hundred sixty-nine is divisible by 3? The first digit is a 3. 9 is divisible by 3. sixty-nine is divisible by 3. 3 plus 6 plus 9 is divisible by 3. @juliannescriv

OpenStudy (anonymous):

b

OpenStudy (anonymous):

3.05]What is the simplified mixed number form of 3.84? three and twenty-one over twenty-five three and forty-two over fifty three and eighty-four over ten three and eighty-four over one hundred @juliannescriv

OpenStudy (anonymous):

a

OpenStudy (anonymous):

[3.02]Which of the following is a pair of equivalent fractions? three-twelfths and nine thirty-sixths two-sevenths and four twenty-eighths five twenty-fifths and six thirty-sixths one-eighth and two-ninths @juliannescriv

OpenStudy (anonymous):

a

OpenStudy (anonymous):

3.04]Helena made trail mix and wants to list the ingredients in order of increasing weight. She used one and five eighths lb of mixed nuts, five eighths lb of chocolate covered raisins, and one and two eighths lb of dried mixed fruit. If she decides to include one and seven eighths lb of sunflower seeds, which list shows the order of ingredients from least weight to greatest weight? seeds, nuts, raisins, fruit nuts, seeds, raisins, fruit raisins, fruit, nuts, seeds raisins, fruit, seeds, nuts @juliannescriv

OpenStudy (anonymous):

a

OpenStudy (anonymous):

@juliannescriv 3.03]If fifteen is the numerator of a proper fraction, which of the following could be the denominator? 1 3 thirteen twenty

OpenStudy (anonymous):

d

OpenStudy (anonymous):

3.04]Which fraction falls between five-eighteenths and five-ninths? one-eighteenth two-ninths four-ninths thirteen-eighteenths @juliannescriv

OpenStudy (anonymous):

c

OpenStudy (anonymous):

3.01]Which of the following is not true? 5 and 7 are the only prime factors of thirty-five. The prime factorization of eighteen is 2 times 9. 2 is the only prime factor of 8. The prime factorization of twelve is 2 squared times 3. @juliannescriv

OpenStudy (anonymous):

b

OpenStudy (anonymous):

3.02]Which pair of numbers has 8 as a common factor? 2 and 4 twenty-four and eighty-six forty-two and seventy-two thirty-two and fifty-six @juliannescriv

OpenStudy (anonymous):

d